# Equivalence of Privacy and Stability with Generalization Guarantees in Quantum Learning

### v1

@@ Appendix A Technical Discussion related to Information-Theoretic Admissibility (ITA)
$\Gamma_{w}\circ\mathcal{N}^{\prime}_{w}(\rho_{s})=\mathcal{N}_{w}(\rho_{s})\quad\forall\,(s,w)\in\mathcal{S}\times\mathcal{W}.$
@@
Appendix B Comparison between Theorem 2 and [15, Proposition 10]
In [15, Proposition 10], the authors derived an upper bound on the Holevo information for quantum $(\varepsilon,\delta)$-LDP quantum channels, as stated in Eq. (209) of [15]. However, one of the authors of [15] later clarified to the authors of the present paper [21] that the phrase “for quantum $(\varepsilon,\delta)$-LDP quantum channels” was a typographical error.
then, the following bound holds:
In fact, if we strengthen our assumption to match theirs—namely, require (9) for any distinct $s,s^{\prime}\in\mathcal{S}$—then part $(i)$ of [7, Corollary 3] recovers the same bound as (16), thereby aligning our result with the corrected version of [15, Proposition 10].
Appendix C Comparison between Theorem 2 and [6, Appendix C.7]
C-A Assumption in [6, Appendix C.7]
Their first part discusses the Holevo information under a certain condition. However, a careful examination of the proof in [6, Appendix C.7] reveals that the argument relies on a stronger assumption than their statement as follows. In this place, the authors claim to prove the following bound on the Holevo information.
for all $0\preceq O\preceq\mathbb{I}^{\mathrm{test;hyp}}$ and $\rho^{\mathrm{test;train}}_{1},\rho^{\mathrm{test;train}}_{2}\in\mathcal{D}(\mathcal{H}^{\mathrm{test;train}})$. In other words, the proof appears to require that $\Lambda^{\mathcal{A}}_{s,w}$, which is locally $\varepsilon$-LDP on $\mathcal{H}^{\mathrm{hyp}}$, also preserves $\varepsilon$-LDP globally when extended to the joint space $\mathcal{H}^{\mathrm{test;hyp}}$. Crucially, (19) does not imply (20). Indeed, by [23, Theorem 4], the identity channel on $\mathcal{H}^{\mathrm{test}}$ fails to satisfy differential privacy for any $\varepsilon\geq 0$, so the composition $\mathbb{I}^{\mathrm{test}}\otimes\Lambda^{\mathcal{A}}_{s,w}$ cannot satisfy $\varepsilon$-LDP solely on the basis of (19). Therefore, there is a gap in the argument of [6, Appendix C.7]: the claimed bound (18) does not follow from their stated assumption (19). That is, one must assume (20) instead of (19). Moreover, the upper bound obtained in (18) involves the term $I(\text{test};\text{train})_{\rho^{\mathcal{A}}{(s,w)}}$. To render this bound meaningful, $I(\text{test};\text{train})_{\rho^{\mathcal{A}}{(s,w)}}$ should also be controlled by some function of the security parameter $\varepsilon$, although our evaluations—such as Theorem 2—do satisfy this requirement.
C-B Security condition in [6, Appendix C.7]
Appendix D Proof of Theorem 1
$\overset{(a)}{\geq}\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\log\mathrm{Tr}\left[e^{\lambda L^{S\bm{\mathfrak{Te}}WB}}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$=\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\log\mathrm{Tr}\left[e^{\lambda\left(L^{S\bm{\mathfrak{Te}}WB}-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]\mathbb{I}^{S\bm{\mathfrak{Te}}WB}\right)}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$~~~-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]$
$\overset{(b)}{\geq}\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\sigma^{S\bm{\mathfrak{Te}}WB}_{\mathcal{N}}\right]-\mathrm{Tr}\left[\lambda L^{S\bm{\mathfrak{Te}}WB}\left(\sigma^{S\bm{\mathfrak{Te}}}\otimes\sigma^{WB^{\prime}}_{\mathcal{N}}\right)\right]-\frac{\lambda^{2}\alpha^{2}}{,}{2}$ (21)
where $(a)$ follows from the variational lower-bound for the quantum relative entropy (see [13, Theorem $5.9$]) and $(b)$ follows from (7) of Definition 5. Further, we can rewrite (21) as follows,
Thus, we have,
This completes the proof of Theorem 1.
Appendix E Proof of Corollary 2
@@
$=e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\sigma)]+g_{k}(\delta),$
@@
Appendix F Proofs of Theorem 2, Corollary 3 and Corollary 4
F-A Proof of Theorem 2
The proof of Theorem 2 above relies on the Claims 1 and 2 below. The proofs of these claims are given in Appendix I and J.
Claim 1.
Claim 2.
Consider $\rho$ and $\sigma$ be two quantum states over Hilbert space $\mathcal{H}$ such that $\rho\ll\sigma$ and $\sigma$ is a finite mixture of probability distributions such that $\sigma=\sum_{b=1}^{m}P(b)\sigma_{b}$, where $\sum_{b=1}^{m}P(b)=1$, and $\rho\ll\sigma_{b}$ for all $b\in[m]$. Then, $D(\rho\|\sigma)\leq\min_{b\in[m]}\left\{D\left(\rho\|\sigma_{b}\right)-\log P(b)\right\}.$
$I[S;WB^{\prime}]$ $\leq D(\sigma^{SB}\|\sigma^{S}\otimes\omega^{B})$
(Step $\bf 1$}) Consider $\omega^{B}$ to be a uniform mixture of $\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}),$ over all the types $\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}$ i.e. $\omega^{B}:=\frac{1}{\left|T^{n}_{\left|\mathcal{Z}\right|}\right|}\sum_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}\mathcal{N}^{\bf f}(\rho_{\mathbf{f}})$, then, using Claim 2 and the fact that $\left|T^{n}_{\left|\mathcal{Z}\right|}\right|\leq(n+1)^{\left|\mathcal{Z}\right|-1}$ (see [12, Eq. $6.18$]), Eq. (25) can be upper-bounded as follows,
$~~~I[S;WB^{\prime}]$
$\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\min_{\mathbf{f}\in T^{n}_{\left|\mathcal{Z}\right|}}\left\{D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\mathbf{f}}))-\log\left|T^{n}_{\left|\mathcal{Z}\right|}\right|^{-1}\right\}$
$\leq(\left|\mathcal{Z}\right|-1)\log(n+1).$
$t\in\mathbb{N}:t\in[1,n].$ (26)
Then, for any $s\in\mathcal{S}$ the following holds,
$\min_{s^{\prime}\in T_{{\bf g}^{(s)}}}d(s,s^{\prime})$ $\leq(\left|\mathcal{Z}\right|-1)\frac{n}{t},$ (27)
where $d(s,s^{\prime})$ is the Hamming distance between $s$ and $s^{\prime}$.
(Step $\bf 3$}) We now prove Theorem 2 using the grid covering technique discussed in the proof of [17, Proposition 2]. Fix $\omega^{B}=\sum_{{\bf f}\in T^{\prime}}\frac{1}{\left|T^{\prime}\right|}\mathcal{N}^{\bf f}(\rho_{\bf f})$, where $T^{\prime}$ is the collection of the center points of all the smaller grids obtained in Step 2. Then, using Claim 2 and the fact that $\left|T^{\prime}\right|\leq t^{\left|\mathcal{Z}\right|-1}$, we have,
$I[S;WB^{\prime}]$ $\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\min_{{\bf f}\in T^{\prime}}\left\{D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{\bf f}(\rho_{\bf f}))+(\left|\mathcal{Z}\right|-1)\log t\right\}$
$\leq\sum_{s\in\mathcal{S}}P^{\otimes n}_{Z}(s)\left(D(\mathcal{N}^{s}(\rho_{s})\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))+(\left|\mathcal{Z}\right|-1)\log t\right).$ (28)
(Step $\bf 4$}) We will now analyze the first term in the RHS of (28) by using Claim 1 and [18, Lemma $6.9$]. Toward this, in [18, Lemma $6.9$], let $\rho=\mathcal{N}^{s}(\rho_{s})$ and $\sigma=\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}})$. Thus, [18, Lemma $6.9$] implies that there exists a quantum state $\mathcal{N}^{s}(\rho_{s})^{\prime}$ in the close vicinity of $\mathcal{N}^{s}(\rho_{s})$ such that $D_{\max}(\mathcal{N}^{s}(\rho_{s})^{\prime}\|\mathcal{N}^{{\bf g}^{(s)}}(\rho_{{\bf g}^{(s)}}))\leq f(\varepsilon,\delta)$, where $f(\cdot,\cdot)$ is some function. Thus, using [18, Lemma $6.9$], Claim 1, Assumption (11) and the extension of privacy constraints of $\mathcal{A}$ under $k$-neighboring inputs, we have the following series of inequalities,
where $\varepsilon^{\prime}:=\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}+\log\frac{1}{1-g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta)}$, and $g_{\frac{n(\left|\mathcal{Z}\right|-1)}{t}}(\varepsilon,\delta)=\frac{e^{\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}}-1}{e^{\varepsilon}-1}\delta$. Thus, using Eqs. (28) and (29) we have,
$\hskip-7.0ptI[S;WB^{\prime}]{\leq}\frac{n(\left|\mathcal{Z}\right|-1)\varepsilon}{t}+(\left|\mathcal{Z}\right|-1)\log t+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta),$ (30)
where $h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta):=\log\frac{1}{1-g_{{n(\left|\mathcal{Z}\right|-1)}}(\delta)}+\frac{2}{m}g_{n(\left|\mathcal{Z}\right|-1)}(\delta)$ (observe that $h_{\left|\mathcal{Z}\right|}(\varepsilon,0)=0$) and the last inequality follows from the fact that the grid size $t\geq 1$.
$t^{\star}=n\varepsilon.$ (31)
$I[S;WB^{\prime}]$ $\leq(|\mathcal{Z}|-1)\left(1+\log\left({n\varepsilon}\right)\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta)$
$=(|\mathcal{Z}|-1)\log\left({ne\varepsilon}\right)+h_{\left|\mathcal{Z}\right|}(\varepsilon,\delta).$
This completes the proof of Theorem 2.
F-B Proof of Corollary 3
F-C Proof of Corollary 4
$I[S;WB^{\prime}]$ $\leq(\left|\mathcal{Z}\right|-1)\log(n+1).$
Appendix G Proof of Lemma 1
@@
Appendix H Numerical Comparison of Generalization Error Bounds
In this appendix, we numerically validate our theoretical results by comparing them against the bounds established in [19]. We utilize the classical-quantum toy example described in [19, Section VI] to demonstrate the tightness of our mutual information-based approach.
For this comparison, we evaluate the following two quantities under the condition that the sub-Gaussianity parameters satisfy $\mu=\tau=\alpha$:
  1. 1.
  2. 2.
    The Separated Bound from [19] ($\mathcal{B}_{\text{SEP}}$): We compare against the bound in [19, Eq. (88), Theorem 1], which separates the classical and quantum contributions. In this specific toy example, the first term of their theorem vanishes, reducing the bound to:
As detailed in Figure 3 above, for comparison, we plot (35) and (36) for the example mentioned in [19, Subsection-VI].
Appendix I Proof of Claim 1
@@
$\log\sigma\succeq\log\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
@@
$\mathrm{Tr}[\rho\log\sigma]\leq-\mathrm{Tr}[\rho\log\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
Adding $\mathrm{Tr}[\rho\log\rho]$ to both sides gives
$\mathrm{Tr}[\rho(\log\rho-\log\sigma)]\leq\mathrm{Tr}[\rho(\log\rho-\log\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
@@
Appendix J Proof of Claim 2
We begin by invoking the operator monotonicity of the function $\log(\cdot)$. Since $\sigma\succeq P(b)\sigma_{b}$ for every $b$, we obtain
$\log\sigma\succeq\log P(b)\,\mathbb{I}+\log\sigma_{b}.$ (37)
@@
$D(\rho\|\sigma)$ $=\mathrm{Tr}\left[\rho(\log\rho-\log\sigma)\right]$
$\leq\mathrm{Tr}\left[\rho(\log\rho-\log P(b)\,\mathbb{I}-\log\sigma_{b})\right]$
$=D(\rho\|\sigma_{b})-\log P(b).$ (38)

### v2

@@ Appendix G Technical Discussion related to Information-Theoretic Admissibility (ITA)
$\Gamma_{w}\circ\mathcal{N}^{\prime}_{w}(\rho_{s})=\mathcal{N}_{w}(\rho_{s}),\quad\forall\,(s,w)\in\mathcal{S}\times\mathcal{W}.$
@@
G-A Implications of ITA: Source-Layer Privacy and Quantum Advantage
Resolving the ITA Conflict: Source-Layer Privacy. The impossibility result in Lemma 1 implies that in the classical domain, if a Data Processor is untrusted and executes an ITA (optimal) algorithm, privacy cannot be preserved by the algorithm itself. Since an ITA algorithm extracts all available information, the output effectively reveals the raw input. Consequently, to preserve privacy in the classical untrusted setting, the burden of protection must shift from the algorithm to the input data itself. This is standardly achieved via Input Perturbation or Local Differential Privacy (LDP), where the Respondent applies a local randomization mechanism $\mathcal{M}$ to generate a noisy version $\tilde{s}=\mathcal{M}(s)$. Even if the Untrusted Processor fully recovers $\tilde{s}$ (as allowed under ITA), the underlying sensitive data $s$ remains protected by the noise added at the source. Thus, indistinguishability is enforced at the source layer, making the specific choice of the processor’s algorithm irrelevant to the privacy guarantee.
Quantum Encoding as Intrinsic Source Noise. This necessity for source-layer protection provides a rigorous motivation for our quantum learning framework. In our model, the encoding map $s\mapsto\rho_{s}$ plays a role conceptually equivalent to classical input perturbation, but with a fundamental physical advantage. In the classical setting, distinct data points $s\neq s^{\prime}$ are perfectly distinguishable unless artificial noise is added. In the quantum setting, however, if the encoded states $\{\rho_{s}\}_{s}$ are non-orthogonal, they are physically indistinguishable with certainty. This non-orthogonality introduces an intrinsic, unavoidable uncertainty—effectively “quantum noise”—that prevents even an adversary with unlimited computational power from perfectly distinguishing $s$ from $s^{\prime}$. Therefore, our framework intrinsically embeds privacy into the physical layer. Even if the Untrusted Data Processor employs an ITA algorithm (i.e., performs the optimal Helstrom measurement to extract maximum information), their ability to infer $s$ is fundamentally limited by the non-orthogonality of the encoded states. This confirms that our security condition is robust: privacy is not contingent on the Processor’s cooperation but is guaranteed by the physical nature of the encoding itself.
Appendix H Proof of Corollary 2
@@
$=e^{k\varepsilon}\mathrm{Tr}[\Lambda\mathcal{N}(\sigma)]+g_{k}(\delta).$
@@
Appendix I Proof of Lemma 1
@@
Appendix J Proof of Claim 1
@@
$\ln\sigma\succeq\ln\rho^{\prime}-D_{\max}(\rho^{\prime}\|\sigma)\,\mathbb{I}.$
@@
$\mathrm{Tr}[\rho\ln\sigma]\leq-\mathrm{Tr}[\rho\ln\rho^{\prime}]+D_{\max}(\rho^{\prime}\|\sigma).$
Adding $\mathrm{Tr}[\rho\ln\rho]$ to both sides gives
$\mathrm{Tr}[\rho(\ln\rho-\ln\sigma)]\leq\mathrm{Tr}[\rho(\ln\rho-\ln\rho^{\prime})]+D_{\max}(\rho^{\prime}\|\sigma),$
@@
Appendix K Proof of Claim 2
We begin by invoking the operator monotonicity of the function $\ln(\cdot)$. Since $\sigma\succeq P(b)\sigma_{b}$ for every $b$, we obtain
$\ln\sigma\succeq\ln P(b)\,\mathbb{I}+\ln\sigma_{b}.$ (55)
@@
$D(\rho\|\sigma)$ $=\mathrm{Tr}\left[\rho(\ln\rho-\ln\sigma)\right]$
$\leq\mathrm{Tr}\left[\rho(\ln\rho-\ln P(b)\,\mathbb{I}-\ln\sigma_{b})\right]$
$=D(\rho\|\sigma_{b})-\ln P(b).$ (56)